how to find the value of x in the following,
01.5(x+2y)-(3x+11y)=4
7x-9y-3(x-4y)=32

02. 3x=7+y
5x=9y+41

Answer by rumessa(13)

02. 3x=7+y
5x=9y+41
sol:
3x=7+y
3x-y=7 equ:1
5x=9y+41
5x-9y=41 equ:2
multipy equ:1 by 9 and then subtract equ:2 from in it
9( 3x-y=7) => 27x-9y=63
1( 5x-9y=41)=>5x-9y=41
- + _
22x=22
x=1